# Large Language Models for Cuffless Blood Pressure Measurement From Wearable Biosignals

### v2

@@ Abstract.
Large language models (LLMs) have captured significant interest from both academia and industry due to their impressive performance across various textual tasks. However, the potential of LLMs to analyze physiological time-series data remains an emerging research field. Particularly, there is a notable gap in the utilization of LLMs for analyzing wearable biosignals to achieve cuffless blood pressure (BP) measurement, which is critical for the management of cardiovascular diseases. This paper presents the first work to explore the capacity of LLMs to perform cuffless BP estimation based on wearable biosignals. We extracted physiological features from electrocardiogram (ECG) and photoplethysmogram (PPG) signals and designed context-enhanced prompts by combining these features with BP domain knowledge and user information. Subsequently, we adapted LLMs to BP estimation tasks through fine-tuning. To evaluate the proposed approach, we conducted assessments of ten advanced LLMs using a comprehensive public dataset of wearable biosignals from 1,272 participants. The experimental results demonstrate that the optimally fine-tuned LLM significantly surpasses conventional task-specific baselines, achieving an estimation error of 0.00 $\pm$ 9.25 mmHg for systolic BP and 1.29 $\pm$ 6.37 mmHg for diastolic BP. Notably, the ablation studies highlight the benefits of our context enhancement strategy, leading to an 8.9% reduction in mean absolute error for systolic BP estimation. This paper pioneers the exploration of LLMs for cuffless BP measurement, providing a potential solution to enhance the accuracy of cuffless BP measurement.
@@ 1. Introduction
Blood pressure (BP) is an important physiological indicator that reflects the pressure at which the heart pumps blood into the blood vessels. Both high blood pressure (hypertension) and low blood pressure (hypotension) can lead to a variety of health problems, such as heart disease, stroke, and kidney disease (potter2009controlling, ). Therefore, accurately monitoring BP is crucial for the prevention and management of various cardiovascular diseases. Traditionally, BP measurement has relied on cuff-based techniques (Figure 1 (a)), such as oscillometry and auscultation, which involve the inflation and deflation of a pneumatic cuff around the upper arm. However, these methods suffer from certain limitations, including discomfort, inconvenience, and intermittent measurements.
@@
In recent years, the emergence of cuffless BP measurement techniques has attracted significant interest (ref_1, ). These techniques utilize wearable devices to capture biosignals(ref_2, ), such as electrocardiogram (ECG) and photoplethysmogram (PPG), and then estimates BP values through advanced computational models (Figure 1 (b)). Compared to the traditional cuff-based method, cuffless BP measurement offers better convenience, unobtrusiveness, and continuous monitoring capabilities, which can bring new possibilities to healthcare and health management. Recent advancements in mechanism-driven pulse transit time models and data-driven machine learning and deep learning models have contributed to progress in cuffless BP measurement (ref_3, ). However, further improvement is necessary as these methods have not fully explored the complex patterns of correlation and interactions between biosignals and BP values.
@@
This paper aims to explore the potential of LLMs for cuffless BP measurement using wearable biosignals. The hypothesis is that, through fine-tuning, LLMs can effectively capture the intricate relationships between biosignals and BP values, thereby improving the accuracy of cuffless BP estimation. To achieve this, we extract rich physiological features from the biosignals and embed them as strings into textual prompt templates. By prompt designing and fine-tuning, we construct a model suitable for the cuffless BP measurement task. To the best of our knowledge, this is the first study to utilize LLMs for cuffless BP measurement. Our contributions include:
@@
  • The release of CBPM-LLaMA [^1] Code repository will be made public upon camera-ready. Anonymous double-blind review., an open-source LLM based on LLaMA3-8B specifically fine-tuned for cuffless BP measurement, making it the first of its kind.
@@ 3.1. Signal Feature Extraction
3.2. Prompt Design and Fine-tuning
@@ 3.2.2. Instruction Tuning
3.2.3. Parameter Efficient Fine-tuning (PEFT)
Parameter Efficient Fine-Tuning (PEFT) techniques, such as Low-Rank Adaptation (LoRA), optimize this process by introducing additional trainable parameters in a compact form. These techniques focus on modifying only a limited subset of parameters or incorporating lightweight additional modules, thereby facilitating efficient adaptation with minimal resource expenditure. Moreover, PEFT methods enable the model’s efficient adaptation to specific tasks without necessitating extensive retraining. This adaptability is crucial for customizing LLMs to applications with limited labeled data, such as BP estimation. In our experiments, we employ LoRA, which trains only a small fraction (0.1%-0.5%) of the total parameters of the pre-trained LLM.
@@ 4.2. Implementation Details
The hardware configuration of our computing system consists of 4 RTX A6000 GPUs with 48 GB of memory each. We experimented with LLMs of various scales, including both general-purpose and domain-specific open-source models. Additional, we fine-tuning these models through instruction tuning and the parameter-efficient LoRA technique. The implementations of these models were obtained from unsloth [^5] https://github.com/unslothai/unsloth and Llama Factory (ref_22, ).
To evaluate the performance more reliably and robustly, we performed a 5-fold cross-validation for each experiment, with the other configurations remaining unchanged. The learning rate for LoRA was set within the range of 1e-4 to 2e-4.
@@ 5.1. Performance Metrics
5.2. Fine-tuning Performance
Table 3 presents the results of the instruction tuning LLMs, fine-tuning with LoRA LLMs and baseline methods for BP estimation. As shown, the LLaMA3-8B model achieved the best performance for SBP estimation, while the Qwen2-7B model demonstrated the best performance for DBP estimation. Additionally, the MedAlpaca-7B model, which was obtained by fine-tuning the LLaMA2-7B with medical domain tasks, outperformed the base LLaMA2-7B model. However, the LLaMA3-8B model exhibited superior performance compared to MedAlpaca-7B, likely due to its incorporation of a higher number of parameters. Furthermore, most of the LLMs outperformed the baseline methods. Specifically, when compared to the best-performing baseline (AdaBoost), the LLaMA3-8B model reduced the ME $\pm$ SDE from 1.39 $\pm$ 9.55 mmHg to 0.00 $\pm$ 9.25 mmHg for SBP estimation, and the Qwen2-7B model reduced the ME $\pm$ SDE from 1.62 $\pm$ 7.04 mmHg to 1.29 $\pm$ 6.37 mmHg for DBP estimation. Figure 4 shows the density Bland-Altman and correlation plots of the estimated values from the LLaMA3-8B model against the reference values across all folds. The Bland-Altman plots indicate that most of the estimated points for SBP and DBP are within or close to the 95% limits of agreement, suggesting a close agreement between the predicted BP values and the reference BP values. The correlation plots show that all correlation coefficients are greater than 0.7, indicating a strong correlation between the predicted and reference BP values. These findings suggest that pre-trained LLMs can be effectively adapted to the task of BP estimation through fine-tuning with wearable BP data.
@@ 5.3. Importance of Different Contexts in Prompt Designing for BP Estimation
5.4. Importance of Training Size in Fine-tuning Performance
To understand the required amount of data for fine-tuning, we conducted experiments on the LLaMA3-8B model with fine-tuning varying down-sampled training sizes: 10%, 20%, 30%, 40%, 50%, 60%, 70%, and 80% of the original dataset. The overall results are presented in Figure 5. Notably, even with only 30% of the original dataset, the fine-tuned model already outperforms the baseline model (AdaBoost trained on 80% of the original data) for both SBP and DBP estimation. This finding provides guidance for fine-tuning with LoRA when data and computing resources are limited. As expected, we observe an increasing trend in performance with more training data.
@@ 6. Conclusion
This paper proposes a novel application of using LLMs for cuffless BP estimation from wearable biosignals. We extracted rich physiological features from wearable ECG and PPG signals and integrated them with BP domain-specific knowledge and user information to design context-enhanced prompt templates. Subsequently, we fine-tuned the LLMs with these prompts to adapt them for the BP estimation task. The results obtained from fine-tuning ten LLMs on a wearable dataset comprising 1,272 subjects demonstrate that the performance of the LLMs surpasses that of baseline models. Additionally, our ablation studies further emphasize the importance of incorporating BP domain knowledge and user information into the prompts. These findings demonstrate the feasibility and potential of LLMs in enhancing the performance of cuffless BP estimation.

### v3

@@ Abstract.
Large language models (LLMs) have captured significant interest from both academia and industry due to their impressive performance across various textual tasks. However, the potential of LLMs to analyze physiological time-series data remains an emerging research field. Particularly, there is a notable gap in the utilization of LLMs for analyzing wearable biosignals to achieve cuffless blood pressure (BP) measurement, which is critical for the management of cardiovascular diseases. This paper presents the first work to explore the capacity of LLMs to perform cuffless BP estimation based on wearable biosignals. We extracted physiological features from electrocardiogram (ECG) and photoplethysmogram (PPG) signals and designed context-enhanced prompts by combining these features with BP domain knowledge and user information. Subsequently, we adapted LLMs to BP estimation tasks through instruction tuning. To evaluate the proposed approach, we conducted assessments of ten advanced LLMs using a comprehensive public dataset of wearable biosignals from 1,272 participants. The experimental results demonstrate that the optimally fine-tuned LLM significantly surpasses conventional task-specific baselines, achieving an estimation error of 0.00 $\pm$ 9.25 mmHg for systolic BP and 1.29 $\pm$ 6.37 mmHg for diastolic BP. Notably, the ablation studies highlight the benefits of our context enhancement strategy, leading to an 8.9% reduction in mean absolute error for systolic BP estimation. This paper pioneers the exploration of LLMs for cuffless BP measurement, providing a potential solution to enhance the accuracy of cuffless BP measurement.
@@ 1. Introduction
Blood pressure (BP) is an important physiological indicator that reflects the pressure at which the heart pumps blood into the blood vessels. Both high blood pressure (hypertension) and low blood pressure (hypotension) can lead to a variety of health problems, such as heart disease, stroke, and kidney disease (potter2009controlling, ). Therefore, accurately monitoring BP is crucial for the prevention and management of various cardiovascular diseases. Traditionally, BP measurement has relied on cuff-based techniques [Figure 1 (a)], such as oscillometry and auscultation, which involve the inflation and deflation of a pneumatic cuff around the upper arm. However, these methods suffer from certain limitations, including discomfort, inconvenience, and intermittent measurements.
@@
In recent years, the emergence of cuffless BP measurement techniques has attracted significant interest (ref_1, ). These techniques utilize wearable devices to capture biosignals (ref_2, ), such as electrocardiogram (ECG) and photoplethysmogram (PPG), and then estimates BP values through advanced computational models [Figure 1 (b)]. Compared to the traditional cuff-based method, cuffless BP measurement offers better convenience, unobtrusiveness, and continuous monitoring capabilities, which can bring new possibilities to healthcare and health management. Recent advancements in mechanism-driven pulse transit time models and data-driven machine learning and deep learning models have contributed to progress in cuffless BP measurement (ref_3, ). However, further improvement is necessary as these methods have not fully explored the complex patterns of correlation and interactions between biosignals and BP values.
@@
This paper aims to explore the potential of LLMs for cuffless BP measurement using wearable biosignals. The hypothesis is that, through instruction tuning, LLMs can effectively capture the intricate relationships between biosignals and BP values, thereby improving the accuracy of cuffless BP estimation. To achieve this, we extract rich physiological features from the biosignals and embed them as strings into textual prompt templates. By prompt designing and instruction tuning, we construct a model suitable for the cuffless BP measurement task. To the best of our knowledge, this is the first study to utilize LLMs for cuffless BP measurement. Our contributions include:
@@
  • The release of CBPM-LLaMA [^1] Code repository will be made public upon camera-ready., an open-source LLM based on LLaMA3-8B specifically fine-tuned for cuffless BP measurement, making it the first of its kind.
@@ 3.1. Signal Feature Extraction
3.2. Prompt Design and Instruction Tuning
@@ 4.2. Implementation Details
The hardware configuration of our computing system consists of 4 RTX A6000 GPUs with 48 GB of memory each. We experimented with LLMs of various scales, including both general-purpose and domain-specific open-source models. Additional, we performed instruction tuning on these models. The implementations were obtained from Llama Factory (ref_22, ).
To evaluate the performance more reliably and robustly, we performed a 5-fold cross-validation for each experiment, with the other configurations remaining unchanged.
@@ 5.1. Performance Metrics
5.2. Instruction Tuning Performance
Table 3 presents the results of the instruction tuning LLMs and baseline methods for BP estimation. As shown, the LLaMA3-8B model achieved the best performance for SBP estimation, while the Qwen2-7B model demonstrated the best performance for DBP estimation. Additionally, the MedAlpaca-7B model, which was obtained by fine-tuning the LLaMA2-7B with medical domain tasks, outperformed the base LLaMA2-7B model. However, the LLaMA3-8B model exhibited superior performance compared to MedAlpaca-7B, likely due to its incorporation of a higher number of parameters. Furthermore, most of the LLMs outperformed the baseline methods. Specifically, when compared to the best-performing baseline (AdaBoost), the LLaMA3-8B model reduced the ME $\pm$ SDE from 1.39 $\pm$ 9.55 mmHg to 0.00 $\pm$ 9.25 mmHg for SBP estimation, and the Qwen2-7B model reduced the ME $\pm$ SDE from 1.62 $\pm$ 7.04 mmHg to 1.29 $\pm$ 6.37 mmHg for DBP estimation. Figure 4 shows the density Bland-Altman and correlation plots of the estimated values from the LLaMA3-8B model against the reference values across all folds. The Bland-Altman plots indicate that most of the estimated points for SBP and DBP are within or close to the 95% limits of agreement, suggesting a close agreement between the predicted BP values and the reference BP values. The correlation plots show that all correlation coefficients are greater than 0.7, indicating a strong correlation between the predicted and reference BP values. These findings suggest that pre-trained LLMs can be effectively adapted to the task of BP estimation through instruction tuning with wearable BP data.
@@ 5.3. Importance of Different Contexts in Prompt Designing for BP Estimation
5.4. Importance of Training Size in Instruction Performance
To understand the required amount of data for instruction tuning, we conducted experiments on the LLaMA3-8B model with instruction tuning varying down-sampled training sizes: 10%, 20%, 30%, 40%, 50%, 60%, 70%, and 80% of the original dataset. The overall results are presented in Figure 5. Notably, even with only 30% of the original dataset, the fine-tuned model already outperforms the baseline model (AdaBoost trained on 80% of the original data) for both SBP and DBP estimation. This finding provides guidance for instruction tuning when data and computing resources are limited. As expected, we observe an increasing trend in performance with more training data.
@@ 6. Conclusion
This paper proposes a novel application of using LLMs for cuffless BP estimation from wearable biosignals. We extracted rich physiological features from wearable ECG and PPG signals and integrated them with BP domain-specific knowledge and user information to design context-enhanced prompt templates. Subsequently, we fine-tuned the LLMs with these prompts to adapt them for the BP estimation task. The results obtained from instruction tuning ten LLMs on a wearable dataset comprising 1,272 subjects demonstrate that the performance of the LLMs surpasses that of baseline models. Additionally, our ablation studies further emphasize the importance of incorporating BP domain knowledge and user information into the prompts. These findings demonstrate the feasibility and potential of LLMs in enhancing the performance of cuffless BP estimation.